OpenStudy (anonymous):

use the rules of differentiation to find the derivative of the function. y=pi/3cos(theta)-cos(theta)

OpenStudy (aum):

\(\large y = \frac{\pi}{3}\cos(\theta)-\cos(\theta)\) ?

OpenStudy (aum):

Is that the function?

OpenStudy (anonymous):

yes

OpenStudy (aum):

Is that the function? Or is the first \(\cos(\theta)\) in the denominator?

OpenStudy (aum):

If that is the function, factor out \(\cos(\theta)\) and then differentiate.

OpenStudy (aum):

\(\large y = \frac{\pi}{3}\cos(\theta)-\cos(\theta) = \cos(\theta)(\pi / 3 - 1)\\ \large \frac{dy}{d\theta} = \frac{d}{d\theta}~\cos(\theta)(\pi / 3 - 1) = (\pi / 3 - 1)* \frac{d}{d\theta}\cos(\theta) = ?\)